simlify this expresstion -9H-6+12h+40=22



Found 2 solutions by mathdumb, checkley71:
Answer by mathdumb(3)

-9H+12h-6+40=22 -9H+12h+34=22 -9H+12h=-12 now is -9H suppose to be capitalized? if not -9h+12h=22, then 3h =22 then h= 22/3 or h= 7 1/3 or h=7.333


Answer by checkley71(8403)

-9H-6+12H+40=22 OR 3H=22+6-40 OR 3H=-12 OR H=-4
PROOF -9*-4-6+12*-4+40=22 or 36-6-48+40=22 or 36+40-6-48=22 or 76-54=22 or 22=22